Circular No. 7555

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SATELLITES OF JUPITER S. S. Sheppard, D. C. Jewitt, Y. R. Fernandez, and G. Magnier, Uni-

versity of Hawaii, have reported the discovery with the 2.2-m reflector of a number of faint candidate satellites of Jupiter (cf. IAUC 7525). Orbit computations by B. G. Marsden, Harvard-Smithsonian Center for Astrophysics, confirm ten of the objects to be new jovian satellites, designated S/2000 J 2–J 11. The first nine, discovered during 2000 Nov. 23–26, have retrograde orbits with elements (for the epoch 2001 Apr. 1) in the ranges 0.14–0.16 AU for a, 0.15–0.53 for e, and 146^{o}–166^{o} for i (i.e., similar to Jupiter VIII, IX, XI, XII, and S/1999 J 1); while the tenth, discovered on Dec. 5, has a = 0.08 AU, e = 0.22, i = 29^{o} (similar to Jupiter VI, VII, X, and XIII). Absolute magnitudes are in the range 14.8–16.1. The last observations for all ten satellites were measured by Sheppard from images obtained during 2000 Dec. 27–2001 Jan. 1 by S. Dahm and A. Evans. Complete details, with observations, orbits, and ephemerides, are given on MPEC 2001-A28 and 2001-A29.

SUPERNOVA 2001B IN IC 391 D. W. Xu and Y. L. Qiu, on behalf of Beijing Astronomical Observa-

tory supernova survey, report the discovery of a supernova found on unfiltered images taken on Jan. 3.61 (mag ∼ 15.5) and 4.57 UT (mag ∼ 15.2). SN 2001B is located at α = 4^{h}57^{m}19^{s}.24, δ = +78^{o}11′16′′.5 (equinox 2000.0), which is 5′′.6 west and 8′′.9 south of the nucleus of IC 391. An unfiltered image of the field taken on 2000 Dec. 24.54 (limiting mag 18.3) showed nothing at this position.

SUPERNOVA 2001C IN PGC 19975 T. Puckett, Mountain Town, GA; and and A. Sehgal, Woodinville, WA,

report the discovery of an apparent supernova (mag 14.7) on an unfiltered CCD frame (limiting mag 20.5) taken with the Puckett Observatory 0.60m automated supernova patrol telescope on Jan. 4.09 UT. SN 2001C is located at α = 6^{h}59^{m}36^{s}.10, δ = +59^{o}31′01′′.6 (equinox 2000.0), which is 14′′.7 east and 5′′.7 south of the center of PGC 19975. The new object was also present and confirmed on CCD frames taken on Jan. 5.04, but it does not appear on unfiltered images taken by Puckett on 2000 Oct. 27–31, and it does not appear on Palomar Sky Survey images taken on 1995 Nov. 11 (limiting mag ∼ 21.0) and 1954 Nov. 3 (limiting mag ∼ 19.7).

2001 January 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green